# Evaluating Memory in LLM Agents via Incremental Multi-Turn Interactions

### v1

@@ 3.3 Different Categories of Memory Agents
Long Context Agents
Modern language models often support extended context windows ranging from 128K to over 1M tokens. A straightforward strategy for memory is to maintain a context buffer of the most recent tokens. For example, in a model with a 128K-token limit, the agent concatenates all incoming chunks until the total exceeds the window size. Once the limit is reached, the earliest chunks are evicted in a FIFO (first-in, first-out) manner. This agent design relies solely on positional recency and assumes the model can attend effectively over the current context window.
RAG Agents
RAG-based agents address context limitations by storing past information in an external memory pool and retrieving relevant content as needed. We consider three RAG variants: (1) Simple RAG Agents: All input chunks are stored as raw text. During inference, a keyword or rule-based string matching mechanism retrieves relevant passages. (2) Embedding-based RAG Agents: Each input chunk is embedded and saved. At query time, the agent embeds the query and performs retrieval using cosine similarity between embeddings. (3) Structure-Augmented RAG Agents: After ingesting all input chunks, the agent constructs a structured representation (e.g., knowledge graph or event timeline). Subsequent queries are answered based on this structured memory.
Agentic Memory Agents
Agentic memory agents extend beyond static memory stores by employing agentic loops—iterative reasoning cycles in which the agent may reformulate questions, perform memory lookups, and update its working memory. These agents are designed to simulate a more human-like process of recalling, verifying, and integrating knowledge.
@@ 3.4 Datasets and Agents Formulation
We standardize all datasets into the format: (chunks), (questions), and (answers), where denotes the -th chunk wrapped to construct a user message with instructions of memorizing the content in a sequential input, and represents a single conversation. Each chunk is accompanied by instructions prompting the agent to memorize its contents. Example prompts are provided in Appendix B.1.
When curating datasets like EventQA and FactConsolidation, we deliberately design scenarios where multiple questions follow a single context. This allows us to probe the model’s memory multiple times with one sequential injection. For example, in LME (S*), five contexts are paired with 300 questions (shown in Table 1). This design choice reflects a key trend: as LLMs support increasingly long context windows and memory agents become more capable of handling extended inputs, evaluation datasets must also scale accordingly. Injecting 1M tokens for just one question is resource-inefficient, whereas associating the same input with many questions provides significantly higher utility.
@@ 4.1 Experimental Setup
The datasets are split into four categories and the statistics of all datasets are also shown in Table 1. The evaluation metrics for all datasets are shown in Table 5 in Appendix A, along with more dataset details. Then for the agents, as described in Section 3.3, we consider three categories of agents: Long-Context Agents, RAG agents and Agentic Memory Agents, where RAG Agents can be further split into Simple RAG Agents, Embedding-based RAG Agents and Structure-Augmented RAG Agents.
For chunk size settings, we choose a chunk size of 512 for the RULER-QA, NIAH-MQ, and LME(S*) tasks in AR, as well as for all tasks in CR. This is mainly because these tasks are composed of long texts synthesized from multiple short texts. For other tasks, we use a chunk size of 4096. Considering computational overhead, we uniformly use a chunk size of 4096 for the Mem0 and Cognee methods. We report the settings of the chunk size in Table 15 in Appendix C.
| AR | TTL | LRU | CR | |||||||
| Agent Type | RULER-QA | NIAH-MQ | Bench-QA | LME(S*) | EventQA | MCC | Recom | Bench-Sum | FactCon-SH | FactCon-MH |
| Long-Context Agents | ||||||||||
| GPT-4o | 61.5 | 25.0 | 55.4 | 32.0 | 77.2 | 87.6 | 12.3 | 32.2 | 60.0 | 5.0 |
| GPT-4o-mini | 53.5 | 22.8 | 44.9 | 30.7 | 59.0 | 82.4 | 15.1 | 28.9 | 45.0 | 5.0 |
| GPT-4.1-mini | 74.5 | 94.8 | 45.8 | 55.7 | 82.6 | 75.6 | 16.7 | 41.9 | 36.0 | 5.0 |
| Gemini-2.0-Flash | 73.0 | 83.8 | 53.2 | 47.0 | 67.2 | 84.0 | 8.7 | 23.9 | 30.0 | 3.0 |
| Claude-3.7-Sonnet | 65.0 | 38.0 | 50.6 | 34.0 | 74.6 | 89.4 | 18.3 | 52.5 | 43.0 | 2.0 |
| GPT-4o-mini | 53.5 | 22.8 | 44.9 | 30.7 | 59.0 | 82.0 | 15.1 | 28.9 | 45.0 | 5.0 |
| Simple RAG Agents | ||||||||||
| BM25 | 61.0 | 100.0 | 45.6 | 45.3 | 74.6 | 75.4 | 13.6 | 20.9 | 56.0 | 3.0 |
| Embedding RAG Agents | ||||||||||
| Contriever | 26.5 | 2.5 | 38.1 | 15.7 | 66.8 | 70.6 | 15.2 | 21.2 | 18.0 | 7.0 |
| Text-Embed-3-Small | 52.0 | 7.2 | 44.4 | 48.3 | 63.0 | 70.0 | 15.3 | 25.7 | 28.0 | 3.0 |
| Text-Embed-3-Large | 49.0 | 19.5 | 50.1 | 52.3 | 70.0 | 72.4 | 16.2 | 21.6 | 28.0 | 4.0 |
| NV-Embed-v2 | 83.0 | 73.5 | 51.4 | 55.0 | 72.8 | 69.4 | 13.5 | 20.7 | 55.0 | 6.0 |
| Structure-Augmented RAG Agents | ||||||||||
| RAPTOR | 33.5 | 15.8 | 31.3 | 34.3 | 45.8 | 59.4 | 12.3 | 13.4 | 14.0 | 1.0 |
| GraphRAG | 47.0 | 38.3 | 35.8 | 35.0 | 34.4 | 39.8 | 9.8 | 0.4 | 14.0 | 2.0 |
| HippoRAG-v2 | 71.0 | 67.5 | 45.7 | 50.7 | 67.6 | 61.4 | 10.2 | 14.6 | 54.0 | 5.0 |
| Mem0 | 28.0 | 4.8 | 22.4 | 36.0 | 37.5 | 3.4 | 10.0 | 0.8 | 18.0 | 2.0 |
| Cognee | 33.5 | 4.0 | 19.7 | 29.3 | 26.8 | 35.4 | 10.1 | 2.3 | 28.0 | 3.0 |
| Agentic Memory Agents | ||||||||||
| Self-RAG | 38.5 | 8.0 | 28.5 | 25.7 | 31.8 | 11.6 | 12.8 | 0.9 | 19.0 | 3.0 |
| MemGPT | 39.5 | 8.8 | 20.8 | 32.0 | 26.2 | 67.6 | 14.0 | 2.5 | 28.0 | 3.0 |
@@ 4.2 Overall Performance Comparison
Table 2 presents the overall performance across different benchmarks. We summarize the key findings as follows: (1) Superiority of RAG methods in Accurate Retrieval Tasks. Most RAG Agents are better than the backbone model “GPT-4o-mini” in the tasks within the Accurate Retrieval Category. This matches our intuition where RAG agents typically excel at extracting a small snippet of text that is crucial for answering the question. (2) Superiority of Long-Context Models in Test-Time Learning and Long-Range Understanding. Long-context models achieve the best performance on TTL and LRU. This highlights a fundamental limitation of RAG methods and commercial memory agents, which still follow an agentic RAG paradigm. These systems retrieve only partial information from the past context, lacking the ability to capture a holistic understanding of the input—let alone perform learning across it. (3) Limitation of All Existing Methods on Conflict Resolution. Although being a well-discussed task in model-editing community [32, 11], resolving conflict poses a significant challenge on memory agents. We observe that all methods fail on the multi-hop situation (with achieving at most 6% accuracy). Only long context agents can achieve fairly reasonable results on single-hop scenarios. In Section 4.4.2, we show that current reasoning models can have much better performance, while it does not change the conclusion that Conflict Resolution still poses a significant challenge on all memory mechanisms. (4) Limited Performance of Commercial Memory Agents. Commercial memory agents such as MemGPT and Mem0 exhibit limited performance across a broad range of benchmarks. This shortfall can be attributed to three primary factors. First, these systems frequently fail to capture and preserve sufficient information when storing inputs into memory. For example, Mem0 depends on extracting factual knowledge from inputs, an approach that inherently discards a substantial portion of the original content. As a result, reconstructing inputs and supporting downstream tasks such as question answering becomes significantly more challenging. While Mem0 has demonstrated relatively strong performance on conversational tasks such as LOCOMO—where information density is comparatively low—it tends to perform poorly on benchmarks containing dense informational content, including RULER and -Bench. For tasks emphasizing Time-to-Live (TTL) and Least Recently Used (LRU) retrieval, these limitations are often even more pronounced. Second, both MemGPT and Mem0 rely on retrieval mechanisms that only access a subset of stored information. In the case of Mem0, retrieval is typically performed a single time, similar to conventional RAG methods, constraining the breadth of information available for reasoning. MemGPT, although adopting a more agentic framework that permits multiple retrieval iterations, does not maintain temporal or structural metadata about the stored content. Consequently, the agent is unable to reconstruct longer documents in their original form, which adversely affects performance on LRU and other tasks requiring structured memory retrieval. Finally, methods such as MemGPT depend heavily on embedding-based retrieval mechanisms, which can be insufficient for fine-grained tasks like NIAH, where locating specific, precise information (“the needle in the haystack”) is essential. These embedding-based approaches often struggle to distinguish subtle contextual nuances that are critical for accurate retrieval in such settings.
4.3 Ablation Study
@@
In this Section, we conduct experiments and result analysis along four dimensions: Input Chunk Size, Retrieval TopK, Validation of Dataset, and Computational Latency. More detailed experimental results are provided in Appendix C.
4.4 Ablation Study on Input Chunk Size
@@
To understand how chunk size impacts performance, particularly for RAG methods and agentic memory agents, we conduct an additional analysis where we vary the chunk size while fixing the number of retrieved chunks to 10. The results are presented in Figure 2. From the figure, we observe the following: (1) In the RULER-QA task, reducing chunk size has little effect on BM25 performance. This is expected, as BM25 relies on term-frequency-based scoring and document-level ranking, and does not inherently benefit from finer-grained segmentation beyond the impact on term distributions. In contrast, embedding-based methods—including MemGPT, which uses text-embedding-3-small as its retriever—consistently perform better with smaller chunks. This suggests that finer segmentation improves the granularity and relevance of retrieved results for models that rely on dense semantic representations. (2) In Bench-Sum, however, smaller chunk sizes lead to worse performance. This task requires the agent to summarize an entire conversation, and smaller chunks correspond to fewer available tokens per retrieval. As a result, the agent has access to less context, which degrades summarization quality. The results suggest that, when resources permit, using smaller chunk sizes and increasing the number of retrieval calls during memory construction can improve performance on Accurate Retrieval (AR) tasks. Finer-grained segmentation enhances the relevance of retrieved information, particularly for embedding-based methods. However, for tasks requiring Long-Range Understanding (LRU), varying the chunk size hurts the performance. This is likely because RAG methods are inherently less suited for tasks that demand integration of information across a large, coherent context.
@@ 4.4.1 Ablation Study on Retrieval TopK
In our experiments, although we report results with the number of retrieved chunks set to 10 in Table 2, we also conducted ablation studies with varying retrieval sizes. A subset of these results is visualized in Figure 3, with the full results provided in Table 11. The results indicate that increasing the number of retrieved chunks generally improves performance across most tasks. It is worth noting that, with a chunk size of 4096 tokens, retrieving 10 chunks already yields an input of approximately 40k tokens. This places significant demands on model capacity. Due to this high token volume, we do not evaluate settings with 20 retrieved chunks.
@@
4.4.2 Validation of Dataset FactConsolidation
As the performance of different models on this dataset remains drastically low, we turn to the stronger reasoning model o4-mini and validate our dataset by checking the performance of o4-mini on a smaller version of this dataset. The results are shown in Table 3.
| 512 | 4096 | |||
| M.C. | Q.E. | M.C. | Q.E. | |
| GPT-4o-mini | 0.09 | 5.2 | 0.07 | 5.1 |
| BM25 | 0.11 | 0.79 | 0.10 | 1.8 |
| Contriever | 7.2 | 0.76 | 1.7 | 2.0 |
| Text-Embed-3-Large | 6.3 | 0.54 | 5.4 | 1.8 |
| NV-Embed-v2 | 93.4 | 0.83 | 42.9 | 1.8 |
| RAPTOR | 151 | 0.51 | 133 | 0.60 |
| GraphRAG | 123 | 9.9 | 90.4 | 9.4 |
| HippoRAG-v2 | 817 | 1.1 | 284 | 2.6 |
| Mem0 | 14644 | 1.2 | 2140 | 1.2 |
| Cognee | 8309 | 33.2 | 962 | 4.5 |
| Self-RAG | 8.4 | 2.0 | 6.7 | 1.7 |
| MemGPT | 413 | 10.6 | 93.3 | 11.4 |
4.4.3 Analysis of Computational Latency
To illustrate the latency of various memory agents in terms of (1) Memory Construction (M.C.); (2) Query Execution (Q.E.), we randomly choose 5 examples from RULER-QA2 and LME (S*), and report the latency of various memory agents. This part of experiments is done on a server with Four NVDIA L40 GPU and AMD EPYC 7713 64-Core CPU. We use the NV-Embed-v2 (7B) as the embedding model in HippoRAG-v2. We show the summarized results in Table 4 and put the full results in Table 12 and 14. From the table, we find that using a smaller chunk size requires significantly more time for memory construction, especially for methods such as HippoRAG-v2, Mem0, Cognee, and MemGPT. Meanwhile, methods such as Mem0, Cognee need extremely high resources when constructing the memory, which may pose challenges in real-world applications.
5 Conclusion and Future Work
In this paper, we introduce MemoryAgentBench, a unified benchmark designed to evaluate memory agents across four essential competencies: accurate retrieval, test-time learning, long-range understanding, and conflict resolution. While prior benchmarks focus largely on skill execution or long-context reasoning, MemoryAgentBench fills a critical gap by assessing how agents store, update, and utilize long-term information across multi-turn interactions. To build this benchmark, we restructure existing datasets and propose two new ones—EventQA and FactConsolidation—tailored to stress specific memory behaviors often overlooked in prior work. We evaluate a wide spectrum of agents, including long-context models, RAG-based systems, and commercial memory agents, under a consistent evaluation protocol. Our results reveal that, despite recent advances, current memory agents still exhibit substantial limitations when faced with tasks requiring dynamic memory updates and long-range consistency. One limitation of our work is that the datasets used in MemoryAgentBench are primarily synthetic, which may not fully reflect the characteristics of real-world user conversations. As future work, we aim to collect and curate more realistic, real-world datasets aligned with our four competencies to further enrich and diversify the benchmark and provide more comprehensive evaluations for memory agents.
Acknowledgment
We thank Kevin Lin for engaging in thoughtful discussions around the overall idea and latency evaluation. His input played a role in shaping the evaluation pipeline for the memory agents.
@@ A.1 Accurate Retrieval (AR)
We use five datasets to evaluate the accurate retrieval capability of memory agents.
(1) RULER-QA
We adopt two QA datasets from [15]. These datasets provide multiple synthetic contexts of varying lengths, ranging from 3K to over 200K tokens. We select 100 questions from the datasets with shorter contexts. For each of these 100 questions, we collect all the documents of them, removed duplicate content, and then shuffled and concatenated them to create new long contexts of 197K or 421K tokens, making sure the new contexts containing the gold passage. Since most answers are short informational entities, such as years, names, or yes/no responses, we use substring exact match (SubEM) as the evaluation metric. SubEM measures whether the predicted answer exactly matches the gold answer as a substring, which is a common standard in question answering systems.
(2) NIAH-MQ
We choose a context with a length of 448K tokens, containing 100 groups and a total of 400 queries. We first check whether these queries appeared in the context. Then, we randomly shuffl the queries and their corresponding numbers evenly to prevent them from clustering together in the long context. The primary evaluation criterion is whether the agent can successfully retrieve the correct numbers. Therefore, we use average recall as the evaluation metric.
(3) Bench-En.QA
We borrow this dataset from [53]. It is a QA task using novels in which character names have been replaced. This makes the content more coherent and closer to real-world scenarios. We use ROUGE F1 for this dataset since answers are mostly entity names.

### v2

@@ 3.2 Different Categories of Memory Agents
(1) Long Context Agents Modern language models often support extended context windows ranging from 128K to over 1M tokens. A straightforward strategy for memory is to maintain a context buffer of the most recent tokens. For example, in a model with a 128K-token limit, the agent concatenates all incoming chunks until the total exceeds the window size. Once the limit is reached, the earliest chunks are evicted in a FIFO (first-in, first-out) manner. This agent design relies solely on positional recency and assumes the model can attend effectively over the current context window. (2) RAG Agents RAG-based agents address context limitations by storing past information in an external memory pool and retrieving relevant content as needed. We consider three RAG variants: Simple RAG Agents: All input chunks are stored as raw text. During inference, a keyword or rule-based string matching mechanism retrieves relevant passages. Embedding-based RAG Agents: Each input chunk is embedded and saved. At query time, the agent embeds the query and performs retrieval using cosine similarity between embeddings. Structure-Augmented RAG Agents: After ingesting all input chunks, the agent constructs a structured representation (e.g., knowledge graph or event timeline). Subsequent queries are answered based on this structured memory. (3) Agentic Memory Agents Agentic memory agents extend beyond static memory stores by employing agentic loops—iterative reasoning cycles in which the agent may reformulate questions, perform memory lookups, and update its working memory. These agents are designed to simulate a more human-like process of recalling, verifying, and integrating knowledge.
@@ 3.3 Datasets and Agents Formulation
We standardize all datasets into the format: (chunks), (questions), and (answers), where denotes the -th chunk wrapped to construct a user message with instructions of memorizing the content in a sequential input, and represents a single conversation. Each chunk is accompanied by instructions prompting the agent to memorize its contents. Example prompts are provided in Appendix C.1. When curating datasets like EventQA and FactConsolidation, we deliberately design scenarios where multiple questions follow a single context. This allows us to probe the model’s memory multiple times with one sequential injection. For example, in LME (S*), five contexts are paired with 300 questions (shown in Table 5 in Appendix A). This design choice reflects a key trend: as LLMs support increasingly long context windows and memory agents become more capable of handling extended inputs, evaluation datasets must also scale accordingly. Injecting 1M tokens for just one question is resource-inefficient, whereas associating the same input with many questions provides significantly higher utility.
@@ 4.1 Experimental Setup
The datasets are split into four categories and the statistics of all datasets are also shown in Table 5. The evaluation metrics for all datasets are shown in Table 1, along with more dataset details. For the agents, as described in Section 3.2, we consider three categories of agents: Long-Context Agents, RAG agents and Agentic Memory Agents, where RAG Agents can be further split into Simple RAG Agents, Embedding-based RAG Agents and Structure-Augmented RAG Agents. We give the detailed introduction of each memory agent in Appendix B. For chunk size settings, we choose a chunk size of 512 for the SH-Doc QA, MH-Doc QA, and LME(S*) tasks in AR, as well as for all tasks in SF. This is mainly because these tasks are composed of long texts synthesized from multiple short texts. For other tasks, we use a chunk size of 4096. Considering computational overhead and API cost, we uniformly use a chunk size of 4096 for the Mem0, Cognee, Zep, and MIRIX. We report the settings of the chunk size in Table 14 in Appendix D.
[…]
@@ 4.2 Overall Performance Comparison
Table 4.1 presents the overall performance across different benchmarks. We summarize the key findings as follows: (1) Superiority of RAG methods in Accurate Retrieval Tasks. Most RAG Agents are better than the backbone model “GPT-4o-mini” in the tasks within the Accurate Retrieval Category. This matches our intuition where RAG agents typically excel at extracting a small snippet of text that is crucial for answering the question. (2) Superiority of Long-Context Models in Test-Time Learning and Long-Range Understanding. Long-context models achieve the best performance on TTL and LRU. This highlights a fundamental limitation of RAG methods and commercial memory agents, which still follow an agentic RAG paradigm. These systems retrieve only partial information from the past context, lacking the ability to capture a holistic understanding of the input—let alone perform learning across it. (3) Limitation of All Existing Methods on Selective Forgetting. Although being a well-discussed task in model-editing community (Mitchell et al., 2022; Fang et al., 2024), forgetting out-of-date memory poses a significant challenge on memory agents. We observe that all methods fail on the multi-hop situation (with achieving at most 7% accuracy). Only long context agents can achieve fairly reasonable results on single-hop scenarios. In Section 4.3.4, we show that current reasoning models can have much better performance, while it does not change the conclusion that Selective Forgetting still poses a significant challenge to all memory mechanisms.
4.3 Analysis and Ablation Study
@@
In this section, we present experiments and analysis along five dimensions: input chunk size, retrieval top-, backbone model, dataset validation, and computational latency. Additional results are provided in the appendix, including context length analysis (Appendix D.4), latency and GPU memory usage comparisons (Appendix D.5, D.6), as well as further details on chunk size and top- ablations (Appendix D.2, D.3).
4.3.1 Ablation Study on Input Chunk Size
@@
To understand how chunk size impacts performance, particularly for RAG methods and agentic memory agents, we conduct an additional analysis where we vary the chunk size while fixing the number of retrieved chunks to 10. The results are presented in Figure 2. From the figure, we observe that when resources permit, using smaller chunk sizes and increasing the number of retrieval calls during memory construction can improve performance on Accurate Retrieval (AR) tasks. Finer-grained segmentation enhances the relevance of retrieved information, particularly for embedding-based methods. However, for tasks requiring Long-Range Understanding (LRU), varying the chunk size hurts the performance. This is likely because RAG methods are inherently less suited for tasks that demand integration of information across a large, coherent context.
@@ 4.3.2 Ablation Study on Retrieval TopK
In our experiments, although we report most results with the number of retrieved chunks set to 10 in Table 4.1, we also conducted ablation studies with varying retrieval sizes. A subset of these results is visualized in Figure 3, with the full results provided in Table 9 in Appendix D. The results indicate that increasing the number of retrieved chunks generally improves performance across most tasks. It is worth noting that, with a chunk size of 4096 tokens, retrieving 10 chunks already yields an input of approximately 40k tokens. This places significant demands on model capacity. Due to this high token volume, we do not evaluate settings with 20 retrieved chunks.
4.3.3 Ablation Study on Backbone Model
To investigate how different backbone models impact the performance of various memory agents, we experimented with three different backbone models and selected four representative methods from both the RAG Agents and Agentic Memory categories. The complete experimental results are presented in Table 3. Our findings show that for RAG Agents, once the backbone is sufficiently strong, it no longer serves as the main performance bottleneck. Compared to the default setup, upgrading to a more powerful model like GPT-4.1-mini yields only marginal improvements. In contrast, the main results in Table 4.1 for the MIRIX method under the Agentic Memory category, using a stronger backbone leads to substantial performance gains. This suggests that future advances in backbone models could further boost the effectiveness of Agentic Memory methods.
[…]
| GPT-4o-mini | 74.6 | 13.6 | 19.0 | 48.0 | 38.8 | |
| BM25 | GPT-4.1-mini | 76.4 | 14.0 | 19.4 | 51.0 | 40.2 |
| Gemini-2.0-Flash | 70.8 | 10.0 | 18.9 | 47.0 | 36.7 | |
| GPT-4o-mini | 63.0 | 15.3 | 17.7 | 28.0 | 31.0 | |
| Text-Embed-3-Small | GPT-4.1-mini | 62.0 | 15.5 | 17.9 | 30.0 | 31.4 |
| Gemini-2.0-Flash | 64.0 | 10.3 | 17.2 | 27.0 | 29.6 | |
| GPT-4o-mini | 34.4 | 9.8 | 0.4 | 14.0 | 14.7 | |
| GraphRAG | GPT-4.1-mini | 39.0 | 10.3 | 1.2 | 16.0 | 16.6 |
| Gemini-2.0-Flash | 36.2 | 7.2 | 0.8 | 13.0 | 14.3 | |
| MIRIX | GPT-4o-mini | 29.8 | 9.8 | 9.9 | 14.0 | 15.9 |
| GPT-4.1-mini | 53.0 (23.2) | 10.3 (0.5) | 18.9 (9.0) | 20.0 (6.0) | 25.6 (9.7) |
4.3.4 Validation of Dataset FactConsolidation
@@
As the performance of different models on this dataset remains drastically low, we turn to the stronger reasoning model o4-mini and validate our dataset by checking the performance of o4-mini on a smaller version of this dataset. The results are shown in Table 4. We found that on the 6K version of the FactCon-SH dataset, both models perform well and are generally able to complete the task effectively. However, their performance drops when the context length increases to 32K. Similarly, on the 6K version of the FactCon-MH dataset, the stronger O4-mini reasoning model achieves a decent score of 80.0, but its performance significantly drops to 14.0 when the context window reaches 32K. This indicates that our dataset is solvable under short-context settings, but current memory agents still lack strong long-range reasoning capabilities, making them unable to handle the task when presented with longer historical inputs.
5 Conclusion
In this paper, we introduce MemoryAgentBench, a unified benchmark designed to evaluate memory agents across four essential competencies: accurate retrieval, test-time learning, long-range understanding, and selective forgetting. While prior benchmarks focus largely on skill execution or long-context question answering, MemoryAgentBench fills a critical gap by assessing how agents store, update, and utilize long-term information across multi-turn interactions. To build this benchmark, we restructure existing datasets and propose two new ones—EventQA and FactConsolidation—tailored to stress specific memory behaviors often overlooked in prior work. We evaluate a wide spectrum of agents, including long-context models, RAG-based systems, and commercial memory agents, under a consistent evaluation protocol. Our results reveal that, despite recent advances, current memory agents still exhibit substantial limitations when faced with tasks requiring dynamic memory updates and long-range consistency. One limitation of our work is that due to budget constraints, so we could only conduct experiments on some relatively representative Memory Agents. As future work, we aim to provide more evaluation results for more memory agents.
@@ A.1 Accurate Retrieval (AR)
A.1.1 Definition of AR
The task of accurately retrieving information has been extensively explored in prior work. In the domain of long-context modeling, the Needle-in-a-Haystack (NIAH) task is widely used to evaluate a model’s ability to locate the specific value based on a given key within a lengthy input. In the RAG setting, this corresponds to document-based QA, where the model must identify and extract relevant snippets from one or more documents to answer a query. These snippets might reside in a single location or be distributed across multiple documents. In this paper, we focus on agentic settings, where the “long context” or “multiple documents” become long-form conversations. We define Accurate Retrieval (AR) as the ability of an agent to identify and retrieve important information that may be dispersed throughout a long dialogue history.
A.1.2 Details on AR datasets
We use four datasets to evaluate the accurate retrieval capability of memory agents.
(1) Document Question Answering
We improved two QA datasets from (Hsieh et al., 2024). These datasets provide multiple synthetic contexts of varying lengths, ranging from 3K to over 200K tokens. We select 100 questions from the datasets with shorter context length. For each of these 100 questions, we collect the context and remove duplicate short documents, and then shuffle and concatenate them to create new long documents of 197K or 421K tokens, making sure the new context containing the gold passages. Since most answers are short informational entities, such as years, names, or yes/no responses, we use substring exact match (SubEM) to calculate the accuracy of QA. SubEM measures whether the predicted answer exactly matches the gold answer as a sub-string, which is a common standard in question answering systems.